News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 15 September 2000 9.30am EDT

During last week the observing schedule was replanned to include the dwarf nova SS Cygni which was accepted as a Target Of Opportunity (TOO) on 10 Sept following an outburst from the source. The target was scheduled to coincide with simultaneous observations by EUVE.

The schedule was re-planned a second time following a high radiation event at 3:15pm EDT on Sept 13 that stopped the daily load and activated the science instrument safing sequence (SCS 107). All actions were nominal and assessment of the dump data confirmed that the trigger was due to high proton counts in the EPHIN P4GM channel exceeding the pre-defined threshold of 300 counts/cm^2/s/sr.

The revised schedule from Sept 12 is shown below. The observations of 1H 0707-495 and U Gem scheduled for this week were not performed due to either the TOO or the radiation event, and will be re-scheduled in an upcoming load.

-------------------------------------------
X1543-62                ACIS-S/HETG  Sep 12 
*TOO SS Cygni           ACIS-S/HETG
Eclipse-6 (69.1 min)
Radiation Belts                      Sep 13
*High Radiation Trigger
*TOO SS Cygni           ACIS-S/HETG  Sep 14 
AR Lac-3                ACIS-S/HETG  Sep 15 
Eclipse-7 (53.6 min)
Radiation Belts
AR Lac-2                ACIS-S/HETG  Sep 16 
4U 1626-67              ACIS-S/HETG
-------------------------------------------

The observatory passed through the fifth and sixth eclipses of the Fall 2000 season with nominal power and thermal performance. The final eclipses for the season are as follows:

Eclipse 7    8:29 pm EDT    15 Sept 2000     53.6 min
Eclipse 8   12:20 pm EDT    18 Sept 2000     26.2 min (short eclipse)


The software patch for the short eclipse will be uplinked on 16 Sept.

The planned observations for the coming week are as follows.

------------------------------------------
Tau Sco                 ACIS-S/HETG  Sep 16 
AR Lac-4                ACIS-S/HETG  Sep 17 
Eclipse-8 (26.2 min)                 Sep 18
Radiation Belts
0509-67.5               ACIS-S/HETG
AR Lac-6                ACIS-S/HETG  Sep 19 
0509-67.5               ACIS-S/HETG
AR Lac-5                ACIS-S/HETG  Sep 20 
Tycho SNR               ACIS-S
Radiation Belts
M31 Transient           ACIS-S       Sep 21 
IC 348                  ACIS-I
Betelgeuse              ACIS-I       Sep 22 
Betelgeuse              ACIS-S
Vela                    ACIS-I
BAL 2359-1241           ACIS-S       Sep 23 
Radiation Belts
N23                     ACIS-S       Sep 24

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports